The natural radioactive decay of 235U is by Alpha emission with a half-life of 704 Million...

  1. The natural radioactive decay of 235U is by Alpha emission with a half-life of 704 Million years. Remember that an Alpha particle consists of two protons and two neutrons. What is the radiogenic daughter isotope produced by this decay?

  • 231U

  • 231Th

  • 235Th

  • 238U

  • 206Pb

2. The half-life for the decay of 222Rn gas is about 3.8 days. If we started with one mole of 222Rn gas, how many days (to the nearest 1/10th of a day) would it take until there was only 1/16 of a mole left?
Answer #1

1)

Let the product be X

Balance mass number on both sides

mass of U = mass of He + mass of X

1*235 = 1*4 + M

M = 231

Balance atomic number on both sides

atomic of U = atomic of He + atomic of X

1*92 = 1*2 + A

A = 90

Atomic number 90 is for element Th

So, the product is

231-Th

Answer: 231-Th

2)

After n half life, ratio of mass:

N/No = (1/2)^n

(1/16) = (1/2)^n

(1/2)^4 = (1/2)^n

n = 4

So, it takes 4 half lives

time taken = number of half lives * half life

= 4*3.8 days

= 15.2 days

Answer: 15.2 days